in rectangle ABCD the length of AC is given by (20x+12)cm and the length of diagonal BD is given by (14x+24)cm.find AC and BD
Answer by jorel555(1290)

Since they are both diagonals (AC is a diagonal):
AC=BD
20x+12=14x+24
6x=12
x=2
20x+12=14x+24=52 cms. each!!!!!!